OpenStudy (anonymous):

HELP!! PLEASE JUST TELL ME HOW TO SOLVE IT! Find the length of the minor axis for the ellipse defined by this equation: ((x+2)^2)/(1)+((y+9)^2)/(64)=1

jimthompson5910 (jim_thompson5910):

((x+2)^2)/(1)+((y+9)^2)/(64)=1 ((x-(-2))^2)/(1^2)+((y-(-9))^2)/(8^2)=1 it is now in the form (x-h)^2/a^2+(y-k)^2/b^2=1 where h = -2 k = -9 a = 1 b = 8 the values of and a and b determine the lengths of the major and minor axis the smaller value a = 1 is the length of the semi-minor axis (since the minor axis is shorter than the major axis) so 2a = 2*1 = 2 is the length of the minor axis

OpenStudy (anonymous):

Thank you!